OpenStudy (anonymous):

tan^2 x + sqr3 tanx = 0 solve please

OpenStudy (anonymous):

tan x = sin x / cos x In each quadrant, tan takes on all the values of one sign with magnitudes ranging from 0 to ∞. It is negative in [π/2, π] and in [3π/2, 2π]. Knowing that tan π/6 = tan 30º = (sin π/6)/(cos π/6) = (1/2)(√3/2) = 1/√3, and that tan(x+π/2) = -1/tan x, you can see that tan(2π/3) = tan(π/6 + π/2) = -√3, and that tan(x+π) = tan(x+π/2+π/2) = -1/tan(x+π/2) = tan x, so tan(5π/3) = tan(2π/3) = -√3 so x = 2π/3, 5π/3, if you're in the [0 to 2π] full circle; x = 2π/3, -π/3, if you're in the [-π to π] full circle Lemme know if this helps, if you need more help @~Chance~

OpenStudy (psymon):

You can just factor this. You would solve it the same way you sould solve \[x ^{2}+\sqrt{3}x \implies x(x+\sqrt{3})\] Just do that, but factor out a tan(x) instead. Then you set both factors equal to 0 and solve for each.

OpenStudy (anonymous):

thanks both of you

OpenStudy (anonymous):

Best responce me please?